Blue Angel, Inc., a private firm in the holiday gift industry, is considering a new project. The company currently has a target debt–equity ratio of .45, but the industry target debt–equity ratio is .40. The industry average beta is 1.20. The market risk premium is 8 percent, and the risk-free rate is 6 percent. Assume all companies in this industry can issue debt at the risk-free rate. The corporate tax rate is 40 percent. The project requires an initial outlay of $680,000 and is expected to result in a $100,000 cash inflow at the end of the first year. The project will be financed at the company’s target debt–equity ratio. Annual cash flows from the project will grow at a constant rate of 6 percent until the end of the fifth year and remain constant forever thereafter.

Calculate the NPV of the project. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Cost of equity is given by

ke = risk free rate + \beta X market risk premium = 0.06 + 1.2 x 0.08 = 0.156

Net cost of debt is given by

kd = (1 - tax rate) x gross cost of debt = (1 - 0.40) x 0.06 = 0.036

target debt / equity = 0.45

\Rightarrow debt = 0.45 x equity \Rightarrow equity / (equity + debt) = equity / (equity + 0.45 x equity) = 0.689655172

\Rightarrowdebt / (equity + debt) = 1- 0.689655172 = 0.310344827

Weighted Average Cost of Capital is given by

WACC = {equity / (equity + debt)} x ke  + {debt / (equity + debt)} x kd

= 0.689655172 x 0.156 + 0.310344827 x 0.036 = 0.11875862

i.e. all cash flows must be discounted at 11.875862 %.

The cash flows are as follows

Time (in years) Cash flow (in $)

- 680000

1 100000
2 100000 x 1.06
3 100000 x 1.062
4 100000 x 1.063
5 100000 x 1.064
6 100000 x 1.064
7 100000 x 1.064
8 100000 x 1.064
9 100000 x 1.064
10 ......
11 ......
.....

Therefore the NPV is

- 680000 + 100000 x 1.11875862-1 + 100000 x 1.06 x 1.11875862-2 + 100000 x 1.062 x 1.11875862-3

+ 100000 x 1.063 x 1.11875862-4 + 100000 x 1.064 x 1.11875862-5 + 100000 x 1.064 x 1.11875862-6

+ 100000 x 1.064 x 1.11875862-7 + ..................... {adding first 5 terms and using the formula for the sum of infinite GP for the remaining sum}

= -349655.1528 + { 100000 x 1.064 x 1.11875862-5 / ( 1 - 1.11875862-1) }

= 328943.11